(1+x)dy/dx-xy=x+x^2


1
Expert's answer
2021-10-06T00:02:13-0400

Answer:-

"y'+P(x)y=Q(x)"


"P(x)=-\frac{x}{x+1}" , "Q(x)=x"


"y'+P(x)y=0"


"\frac{dy}{y}=-P(x)dx"


"\int\frac{dy}{y}=-\int P(x)dx"


"ln|y|=-\int P(x)dx"


"y=\pm e^{-\int P(x)dx}"


"\int P(x)dx=\int(-\frac{x}{x+1})dx=-x+ln(x+1)+c"


"y=\frac{ce^x}{x+1}"


"y=\frac{c(x)e^x}{x+1}"


"\frac{d}{dx}c(x)=Q(x)e^{\int P(x)dx}"


"\frac{d}{dx}c(x)=(x^2+x)e^{-x}"


"\int (x^2+x)e^{-x}dx=(-x^2-3x-3)e^{-x}+c"


"y(x)=\frac{e^x((-x^2-3x-3)e^{-x}+c)}{x+1}"